Using standard thermodynamic data from Appendix D, calculate the equilibrium constant at 307 K for the following chemical equilibrium:

FeO(s) + CO(g) ⇄Fe(s) + CO2(g)
Answer #1

FeO(s) + CO(g) ⇄Fe(s) + CO2(g)

G0rxn = G0f products - G0f reactants

              = -394.359 + 0 -(-251.46 -137.168)

               = -5.731Kj/mole

              = -5731J/mole

G0rxn   = -RTlnK

-5731       = -8.314*307lnK

lnK      = -5731/(-8.314*307)

lnK    = 2.2453

K        = 9.443 >>>>>answer